OpenStudy (anonymous):

Lastly, you decide to keep track of your loan four times a month instead of monthly. Solve for the adjusted interest rate. Remember to use the formula A(t)=P[(1+rn)1c]cnt where c = 4. When solving for the adjusted interest rate, be sure to set it equal to 1+rn.

OpenStudy (anonymous):

@Hero

hero (hero):

What are the given values besides c = 4?

OpenStudy (anonymous):

You are going to purchase a new car, but being a responsible consumer means doing a little bit of research first. First, you find the vehicle you are purchasing and its price. Vehicle: Chevy Volt Price: $39,145 Current interest rate: 3% t=-0.2242

OpenStudy (anonymous):

@Hero

hero (hero):

What happened to the 10,000?

OpenStudy (anonymous):

I didn't say in the question i think because i found t

hero (hero):

Basically, this question is the same as the previous question, except we have an extra variable, c. Correct?

OpenStudy (anonymous):

Yes

hero (hero):

So we have the formula \(A(t) = P\left(1 + \dfrac{r}{n}\right)^{nt}\) but we have to add the c variable to it. I'm not exactly sure where the c variable goes. I assume this: \(A(t) = P\left(1 + \dfrac{r}{cn}\right)^{cnt}\)

OpenStudy (anonymous):

They gave me a new variable \[A(t)=P[(1+r/n)^1/c]^cnt\]

hero (hero):

Do you think you might be able to use the draw feature to write out the formula?